# Federico Ceola, Istituto Italiano di Tecnologia, Via San Quirico, 19D 16163 Genova

### v2

@@ Abstract
keywords:
Long-Horizon Robotic Manipulation, Real Robot Dataset, Language Guided Tasks
1 Introduction
@@
Long-horizon manipulation tasks bring forth the importance of perceptual skills in robotic systems. While requiring to execute low-level control policies to solve the sub-tasks in which they are decomposed, these tasks also require high-level planning and reasoning capabilities. Furthermore, as robots become increasingly integrated into human-centered settings, they will be required to understand and follow natural language instructions for such tasks. Understanding natural language instructions offers the opportunity to develop robotic systems that are capable of autonomously learning, generalizing and adapting to novel settings and environments, rather than being explicitly programmed for each task Shridhar et al. (2023); Zitkovich et al. (2023).
@@
The wide availability of datasets for short-horizon manipulation tasks fostered the development of effective learning-based approaches for such tasks Levine et al. (2018); Kalashnikov et al. (2018). In contrast, the robotics literature lacks real-world datasets for long-horizon tasks. With LHManip, our objective is to address this existing gap in the robotics literature, providing data to develop novel approaches to solve real-world long-horizon manipulation tasks, benchmark existing methods for such tasks evaluated only in simulation Yang et al. (2022) on real data, and evaluate generalization properties of state-of-the-art approaches Chen et al. (2023). Our dataset consists of $2020$ tabletop manipulation tasks involving $3333$ everyday objects. For each of these tasks, we provide a natural language description and $1010$ different demonstrations collected via teleoperation. Different demonstrations of the same tasks either involve manipulation of different object instances (e.g. objects with different texture or size) or consider different environment conditions (e.g. different distractors on the table or different initial configurations of the objects involved in the tasks). For each demonstration, we provide visual RGB and depth observations from a wrist-mounted and two static cameras, and robot proprioceptive information. Furthermore, for each timestep of the episode, we provide the cartesian displacement of the end-effector of the robot and the position offset applied to the gripper. Fig. 1 shows a visual observation in LHManip from one of the two external static cameras. This dataset forms part of the larger effort by the Open X-Embodiment collaboration project Open X-Embodiment Collaboration et al. (2023), and this dataset paper serves to provide full details of our contribution.
Figure 3: Sub-tasks decomposition of a sequence from the Place the bowl on the plate and the cup in the bowl matching the color task.
@@ 2 Related Work
Long-horizon manipulation tasks are challenging robotics problems because they require both high-level reasoning capabilities to decompose the tasks in sequences of sub-tasks and low-level reasoning to solve the short-horizon subtasks. Existing approaches based on Hierarchical Reinforcement Learning Nachum et al. (2018); Yang et al. (2022) have been extensively studied to solve long-horizon tasks, but their application in real robotic tasks is hampered by the need of huge amounts of real training data. Other approaches solve long-horizon tasks with methods based on Imitation Learning Mandlekar et al. (2020), combinations of Task and Motion Planning (TAMP) and Reinforcement Learning (RLCheng and Xu (2023), or skills learning with Large Language Models (LLMZhang et al. (2023a), and would benefit from the availability of real-world data for training.
@@
While it has been shown in the literature the benefit of using real robotic datasets to learn short-horizon manipulation tasks such as grasping Levine et al. (2018), pick and place Brohan et al. (2022), pouring and scooping Zhou et al. (2023), object pushing Chi et al. (2023), or combinations of different tasks Walke et al. (2023), datasets for long-horizon tasks are mainly provided in simulated environments. For example, the IKEA furniture assembly environment Lee et al. (2021) provides simulated environments for assembly tasks that require long-horizon manipulation skills. CALVIN Mees et al. (2022), instead, provides a dataset and a benchmark for language-guided long-horizon tasks with the aim of evaluating robot capabilities of learning new skills. This differs from the ALFRED Shridhar et al. (2020) benchmark, that combines seven predefined skills described through language instructions for navigation and manipulation tasks. LoHoRavens Zhang et al. (2023b), instead, is a simulated benchmark composed of ten long-horizon language-conditioned tasks that require at least five pick-and-place steps to be achieved.
@@
To overcome the limitations of learning robotic tasks in simulated environments, FurnitureBench Heo et al. (2023) proposes a benchmark for long-horizon assembly tasks, such as lamp screwing and assembling table legs. While providing a huge amount of real teleoperated data ($51005100$ demonstrations), the considered environments are constrained to tabletop settings with 3D-printed objects with multiple markers attached on them. The dataset used to benchmark results in Rosete-Beas et al. (2022), instead, composes short-horizon language-guided tasks demonstrations into long-horizon tasks in constrained tabletop environments. With our dataset, we aim at overcoming these limitations and to provide a language-grounded dataset to perform cluttered tabletop manipulation tasks on everyday objects, relying only on a single language description for task description, multiple RGB-D observations of the environment and the proprioceptive state of the robot.
3 LHManip
3.1 Experimental Set-Up and Data Collection
We perform the tasks on a Franka Panda 7-DoF arm Haddadin et al. (2022) mounted on a LD-60 Omron mobile base [^1] https://www.ia.omron.com/products/family/3664/dimension.html, keeping the mobile base fixed throughout data collection.
3.2 Dataset
@@
Table 1: Items considered in the LHManip dataset.
@@
3.2.1 Tasks
Task Success Condition
Clean the pan.
The robot picks the sponge and performs a movement on the pan.
Cook the capsicum and place it on a plate.
The robot picks the capsicum, puts it in the pan, and then places it on a plate.
Cook the vegetables.
All the vegetables are in the pan.
Dry the plate.
The robot grasps a tissue and performs a movement on the plate.
Hide the teddy bear in the red bowl.
The teddy bear is in the red bowl, and there is an object on it.
Match the cups with the appropriate bowls.
The cups are in the bowls of the same color.
Place the bowl on the plate and the cup in the bowl matching the color.
The bowl is on the plate of the same color, and the cup is in the bowl of the same color.
Place the bowls on the appropriate plates.
The bowls are on the plates of the same color.
Prepare two cups of tea.
The two tea bags are placed in two different cups.
Put a highlighter on each book.
There is a highlighter on each book.
Put the ball in the red pot.
The ball is in the red pot.
Roll the dices in the bowl.
The dices are thrown in the bowl.
Serve the vegetables in different plates.
The vegetables in the pan are moved to different plates.
Set the table.
The fork and the spoon are to the right of the plate, and the cup is beyond the plate from the robot’s point of view.
Sort the balls from left to right in order of size.
The balls are sorted in decreasing order from left to right in the robot’s point of view.
Stack green blocks.
The three green blocks are stacked.
Stack the bowls.
The three bowls on the table are stacked.
Stack the cups.
The three cups on the table are stacked.
Throw away the rubbish paper.
The rubbish paper is in the trash bin.
Water the potted plant and put the can on the plate.
The robot performs a movement with the can close to the potted plant and then puts the can on the plate.
Table 2: Tasks Overview. We provide the list of tasks in the dataset, specifying the conditions that must be met to consider the task execution successful.
We report in Tab. 2 an overview of the $2020$ tasks considered in our dataset. The tasks are performed in a highly-cluttered tabletop environment and require the robot to manipulate everyday objects. We provide the list of objects used in the dataset in Tab. 1. The considered tasks are composed by at least two sub-tasks that the robot must complete to successfully achieve the task. We report in Fig. 3 an exemplar sequence from the Place the bowl on the plate and the cup in the bowl matching the color task. As it can be noticed, to achieve the task, the robot needs to perform several sub-tasks: it firstly grasps the orange bowl and places it on the orange plate. Then it picks the orange cup and places it in the orange bowl.
@@ 3.2.1 Tasks
Furthermore, with our dataset we aim at providing data to solve tasks that require high-level reasoning capabilities. In that we specify the task at hand via a natural language instruction, without providing any other low-level environment descriptions (i.e. object details like shape, spatial information, or color). We believe that the robot must be able to infer this information autonomously from the environment. For these reasons, for each task, we provide different variations of either the initial placement of the objects in the environments or variations of the objects involved to achieve the task. Fig. 4 shows two exemplar task variations for two different tasks, where the robot is required to solve the same task with different objects.
3.2.2 Observation and Action Space
@@
LHManip provides a set of visual and proprioceptive information that captures the robot movement and the action that the robot is required to perform via teleoperation. In the dataset, we provide for each timestep observations and actions as reported in Tab. 3. We control the robot in a non-blocking mode at $3030$ Hz.
@@
While we control the cartesian end-effector of the robot and the aperture of the gripper, in the dataset, we also provide joint-level information such as joint positions and velocities.
@@
Observation Description Details
Main static RGB camera
$640\times 480\times 3640 × 480 × 3$
Main static Depth camera
$848\times 480\times 1848 × 480 × 1$
Cameras
Secondary static RGB camera
$640\times 480\times 3640 × 480 × 3$
Secondary static Depth camera
$848\times 480\times 1848 × 480 × 1$
Wrist-mounted RGB camera
$640\times 480\times 3640 × 480 × 3$
Wrist-mounted Depth camera
$848\times 480\times 1848 × 480 × 1$
End-effector position
(x, y, z) w.r.t. root_frame
End-effector orientation
(x, y, z, w) quaternion w.r.t. root_frame
Proprioceptive
Robot joint angles
$77$ values in rad
Gripper position
$22$ values in [0, 0.0404]
Robot joint velocities
$77$ values in rad/s
Instruction
Natural language instruction
String
Action Description Details
End-effector position displacement
(x, y, z) w.r.t. root_frame
Robot Action
End-effector orientation displacement
(x, y, z, w) quaternion w.r.t. root_frame
Gripper opening displacement
$11$ value in [-0.0808, 0.0808]
Table 3: Observations and Actions provided in LHManip.
3.2.3 Dataset Access
The LHManip dataset is publicly available [^3] https://www.dropbox.com/scl/fi/6t717h5mo5kyhq521qavb/long_horizon_manipulation_dataset.zip?rlkey=rk4wsxp464x5bt4a8tgz563ne&dl=0 and can be downloaded as a single .zip file. We provide data in .png format for RGB and Depth images, and in .pkl files for numerical and textual information. We provide a Python snippet code and the instructions to parse the dataset on GitHub [^4] https://github.com/fedeceola/LHManip. Furthermore, LHManip is included as part of the larger dataset released in the Open X-Embodiment project Open X-Embodiment Collaboration et al. (2023). We release additional depth and the unprocessed sensory data together with this white paper, as well as the code to preprocess this dataset [^5] https://github.com/fedeceola/rlds_dataset_builder back into the desired RLDS [^6] https://github.com/google-research/rlds data format as required by the Open X-Embodiment project.
4 Conclusion
@@
In the perspective of developing robots that can interact with humans in everyday environments, we believe that they must possess the ability to address long-horizon tasks based on a single high-level natural instruction, relying solely on visual and proprioceptive feedback. Existing real datasets are either constrained to simplified environments or have long-horizon instructions composed of short-horizon task descriptions. LHManip, instead, presents several challenges, such as natural language instruction understanding, visual perception of the environment in the presence of changing challenging conditions, and learning of low-level control policies for sub-tasks execution.

### v3

# LHManip: A Dataset for Long-Horizon Language-Grounded Manipulation Tasks in Cluttered Tabletop Environments

@@ Abstract
I Introduction
@@
Long-horizon manipulation tasks bring forth the importance of perceptual skills in robotic systems. While requiring to execute low-level control policies to solve the sub-tasks in which they are decomposed, these tasks also require high-level planning and reasoning capabilities. Furthermore, as robots become increasingly integrated into human-centered settings, they will be required to understand and follow natural language instructions for such tasks. Understanding natural language instructions offers the opportunity to develop robotic systems that are capable of autonomously learning, generalizing and adapting to novel settings and environments, rather than being explicitly programmed for each task [16, 22].
@@
The wide availability of datasets for short-horizon manipulation tasks fostered the development of effective learning-based approaches for such tasks [9, 7]. In contrast, the robotics literature lacks real-world datasets for long-horizon tasks. With LHManip, our objective is to address this existing gap in the robotics literature, providing data to develop novel approaches to solve real-world long-horizon manipulation tasks, benchmark existing methods for such tasks evaluated only in simulation [18] on real data, and evaluate generalization properties of state-of-the-art approaches [2]. Our dataset consists of $2020$ tabletop manipulation tasks involving $3333$ everyday objects. For each of these tasks, we provide a natural language description and $1010$ different demonstrations collected via teleoperation. Different demonstrations of the same task either involve manipulation of different object instances (e.g. objects with different texture or size) or consider different environment conditions (e.g. different distractors on the table or different initial configurations of the objects involved in the tasks). For each demonstration, we provide visual RGB and depth observations from a wrist-mounted and two static cameras, and robot proprioceptive information. Furthermore, for each timestep of the episode, we provide the cartesian displacement of the end-effector of the robot and the position offset applied to the gripper. Fig. 1 shows a visual observation in LHManip from one of the two external static cameras. This dataset forms part of the larger effort by the Open X-Embodiment collaboration project [13], and this dataset paper serves to provide full details of our contribution.
Figure 3: Sub-tasks decomposition of a Place the bowl on the plate and the cup in the bowl matching the color sequence.
@@ II Related Work
Long-horizon manipulation tasks are challenging robotics problems because they require both high-level reasoning capabilities to decompose the tasks in sequences of sub-tasks and low-level reasoning to solve the short-horizon subtasks. Existing approaches based on Hierarchical Reinforcement Learning [12, 18] have been extensively studied to solve long-horizon tasks, but their application in real robotic tasks is hampered by the need of huge amounts of real training data. Other approaches solve long-horizon tasks with methods based on Imitation Learning [10], combinations of Task and Motion Planning (TAMP) and Reinforcement Learning (RL[3], or skills learning with Large Language Models (LLM[19], and would benefit from the availability of real-world data for training.
Task Success Condition
Clean the pan. The robot picks the sponge and performs a movement on the pan.
Cook the capsicum and place it on a plate. The robot picks the capsicum, puts it in the pan, and then places it on a plate.
Cook the vegetables. All the vegetables are in the pan.
Dry the plate. The robot grasps a tissue and performs a movement on the plate.
Hide the teddy bear in the red bowl. The teddy bear is in the red bowl, and there is an object on it.
Match the cups with the appropriate bowls. The cups are in the bowls of the same color.
Place the bowl on the plate and the cup in the bowl matching the color. The bowl is on the plate of the same color, and the cup is in the bowl of the same color.
Place the bowls on the appropriate plates. The bowls are on the plates of the same color.
Prepare two cups of tea. The two tea bags are placed in two different cups.
Put a highlighter on each book. There is a highlighter on each book.
Put the ball in the red pot. The ball is in the red pot.
Roll the dices in the bowl. The dices are thrown in the bowl.
Serve the vegetables in different plates. The vegetables in the pan are moved to different plates.
Set the table. The fork and the spoon are to the right of the plate, and the cup is beyond the plate from the robot’s point of view.
Sort the balls from left to right in order of size. The balls are sorted in decreasing order from left to right in the robot’s point of view.
Stack green blocks. The three green blocks are stacked.
Stack the bowls. The three bowls on the table are stacked.
Stack the cups. The three cups on the table are stacked.
Throw away the rubbish paper. The rubbish paper is in the trash bin.
Water the potted plant and put the can on the plate. The robot performs a movement with the can close to the potted plant and then puts the can on the plate.
TABLE I: Tasks. We provide the list of tasks in the dataset, specifying the conditions that must be met to achieve them.
@@
While it has been shown in the literature the benefit of using real robotic datasets to learn short-horizon manipulation tasks such as grasping [9], pick and place [1], pouring and scooping [21], object pushing [4], or combinations of different tasks [17], datasets for long-horizon tasks are mainly provided in simulated environments. For example, the IKEA furniture assembly environment [8] provides simulated environments for assembly tasks that require long-horizon manipulation skills. CALVIN [11], instead, provides a dataset and a benchmark for language-guided long-horizon tasks with the aim of evaluating robot capabilities of learning new skills. This differs from the ALFRED [15] benchmark, that combines seven predefined skills described through language instructions for navigation and manipulation tasks. LoHoRavens [20], instead, is a simulated benchmark composed of ten long-horizon language-conditioned tasks that require at least five pick-and-place steps to be achieved.
@@
To overcome the limitations of learning robotic tasks in simulated environments, FurnitureBench [6] proposes a benchmark for long-horizon assembly tasks, such as lamp screwing and assembling table legs. While providing a huge amount of real teleoperated data ($51005100$ demonstrations), the considered environments are constrained to tabletop settings with 3D-printed objects with multiple markers attached on them. The dataset used to benchmark results in [14], instead, composes short-horizon language-guided tasks demonstrations into long-horizon tasks in constrained tabletop environments. With our dataset, we aim at overcoming these limitations and to provide a language-grounded dataset to perform cluttered tabletop manipulation tasks on everyday objects, relying only on a single language description for task description, multiple RGB-D observations of the environment and the proprioceptive state of the robot.
III LHManip
@@
TABLE II: Items considered in the LHManip dataset.
Observation Description Details
Main static RGB camera $640\times 480\times 3640 × 480 × 3$
Main static Depth camera $848\times 480\times 1848 × 480 × 1$
Cameras Secondary static RGB camera $640\times 480\times 3640 × 480 × 3$
Secondary static Depth camera $848\times 480\times 1848 × 480 × 1$
Wrist-mounted RGB camera $640\times 480\times 3640 × 480 × 3$
Wrist-mounted Depth camera $848\times 480\times 1848 × 480 × 1$
End-effector position (x, y, z) w.r.t. root_frame
End-effector orientation (x, y, z, w) quaternion w.r.t. root_frame
Proprioceptive Robot joint angles $77$ values in rad
Gripper position $22$ values in [0, 0.0404]
Robot joint velocities $77$ values in rad/s
Instruction Natural language instruction String
Action Description Details
End-effector position displacement (x, y, z) w.r.t. root_frame
Robot Action End-effector orientation displacement (x, y, z, w) quaternion w.r.t. root_frame
Gripper opening displacement $11$ value in [-0.0808, 0.0808]
TABLE III: Observations and Actions provided in LHManip.
III-A Experimental Set-Up and Data Collection
We perform the tasks on a Franka Panda 7-DoF arm [5] mounted on a LD-60 Omron mobile base [^1] https://www.ia.omron.com/products/family/3664/dimension.html, keeping the mobile base fixed throughout data collection.
III-B Dataset
@@
III-B1 Tasks
@@
We report in Tab. I an overview of the $2020$ tasks considered in our dataset. The tasks are performed in a highly-cluttered tabletop environment and require the robot to manipulate everyday objects. We provide the list of objects used in the dataset in Tab. II. The considered tasks are composed by at least two sub-tasks that the robot must complete to successfully achieve the task. We report in Fig. 3 an exemplar sequence from the Place the bowl on the plate and the cup in the bowl matching the color task. As it can be noticed, to achieve the task, the robot needs to perform several sub-tasks: it firstly grasps the orange bowl and places it on the orange plate. Then it picks the orange cup and places it in the orange bowl. Furthermore, with our dataset we aim at providing data to solve tasks that require high-level reasoning capabilities. In that we specify the task at hand via a natural language instruction, without providing any other low-level environment descriptions (i.e. object details like shape, spatial information, or color). We believe that the robot must be able to infer this information autonomously from the environment. For these reasons, for each task, we provide different variations of either the initial placement of the objects in the environment or variations of the objects involved to achieve the task. Fig. 4 shows two exemplar task variations for two different tasks, where the robot is required to solve the same task with different objects.
III-B2 Observation and Action Space
@@
LHManip provides a set of visual and proprioceptive information that captures the robot movement and the action that the robot is required to perform via teleoperation. In the dataset, we provide for each timestep observations and actions as reported in Tab. III. We control the robot in a non-blocking mode at $3030$ Hz.
@@
While we control the cartesian end-effector of the robot and the aperture of the gripper, in the dataset, we also provide joint-level information such as joint positions and velocities which can be used for keypoints extraction and enable task training with keypoint-based methods as PerAct [16].
@@
III-B3 Dataset Access
The LHManip dataset is publicly available [^3] https://huggingface.co/datasets/fedeceola/LHManip. We provide data in .png format for RGB and Depth images, and in .pkl files for numerical and textual information. We provide a Python snippet code and the instructions to parse the dataset on GitHub [^4] https://github.com/fedeceola/LHManip. Furthermore, LHManip is included as part of the larger dataset released in the Open X-Embodiment project [13]. We release additional depth and unprocessed sensory data together with this white paper, as well as the code to preprocess this dataset [^5] https://github.com/fedeceola/rlds_dataset_builder back into the desired RLDS [^6] https://github.com/google-research/rlds data format as required by the Open X-Embodiment project.
IV Conclusion
@@
In the perspective of developing robots that can interact with humans in everyday environments, we believe that they must possess the ability to address long-horizon tasks based on a single high-level natural instruction, relying solely on visual and proprioceptive feedback. Existing real datasets are either constrained to simplified environments or have long-horizon instructions composed of short-horizon task descriptions. LHManip, instead, presents several challenges, such as natural language instruction understanding, visual perception of the environment in the presence of changing challenging conditions, and learning of low-level control policies or keypoint-based methods for sub-tasks execution.